OpenStudy (anonymous):

solve the equation -4(x+10)-6=-3(x-2)

OpenStudy (anonymous):

you need to distribute the -4 and the -3 so: -4(x + 10) -6 = -4x -40 -6 -3(x -2) = -3x + 6 -4x -40 -6 = -3x + 6 add the -40 + -6 -4x -46 = -3x + 6 plus 4x to both sides -46 = x + 6 minus 6 from both sides to solve for x x = ?

OpenStudy (anonymous):

-52?

OpenStudy (anonymous):

yes! :) good job

OpenStudy (anonymous):

thanks